OpenStudy (anonymous):

find the indefinite integral of cos^3(12x)

OpenStudy (anonymous):

@wio can you help me please?

OpenStudy (anonymous):

whta's the range?

OpenStudy (anonymous):

Not provided.

OpenStudy (anonymous):

idk then, my friend is good at this, I forgot this stuff..

OpenStudy (anonymous):

can you take him in?

OpenStudy (anonymous):

tag him in this question I mean

OpenStudy (ipwnbunnies):

I'll give a tip and say this is a tedious integration by parts problem I think.

OpenStudy (anonymous):

Looool, thanks anyways for your tip :).

OpenStudy (anonymous):

@surjithayer try this

OpenStudy (fibonaccichick666):

Wait it's easy, if I gave you \[\int cos^2(12x)cos(12x) dx\] That help

OpenStudy (anonymous):

and then integration by parts?

OpenStudy (fibonaccichick666):

Not quite what do you know cos^2 is equal to?

OpenStudy (anonymous):

\[\cos 3x=4\cos ^3x-3\cos x\] \[\cos ^3x=\frac{ 1 }{4 }\left( \cos 3x+3\cos x \right)\]

OpenStudy (anonymous):

I know that cos^2(x) = 1- sin^2(x)

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (fibonaccichick666):

Yup

OpenStudy (ipwnbunnies):

Wow, that actually makes it easier. Do wut Chick said!

OpenStudy (fibonaccichick666):

So plug that in place then just do u sub no parts needed

OpenStudy (ipwnbunnies):

It's...It's almost like cheatin...In math. :o

OpenStudy (anonymous):

\[\cos ^3\left( 12x \right)=\frac{ 1 }{4 }\left\{ \cos \left( 36x \right)+3\cos \left( 12x \right) \right\}\]

OpenStudy (anonymous):

so cos^2(12x)(cos(x) = (1-sin^2(12x))(cos(x)? and then let u = cos x?

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (fibonaccichick666):

Distribute first

OpenStudy (anonymous):

can you solve it please? Thank you

OpenStudy (fibonaccichick666):

Then separate the integrals and yes cos for u

OpenStudy (fibonaccichick666):

Sorry sin for u not cos

OpenStudy (anonymous):

how do I treat the 12 that is in the sine function?

OpenStudy (fibonaccichick666):

Just another sub make it w or something

OpenStudy (fibonaccichick666):

If you show me your steps, it's easier to assist

OpenStudy (anonymous):

oh okay! Thanks I can take it on from here. It was the 12x that was ruining my day.

OpenStudy (fibonaccichick666):

oh yea, just a basic whats its deriv, then pull out the constant

OpenStudy (fibonaccichick666):

np and avoid integration by parts when possible, it's long and annoying

OpenStudy (anonymous):

I will divide by 12 and integrate it as if the inside was an x. Thanks.

OpenStudy (anonymous):

But I wanna ask, could this be done using integration by parts?

OpenStudy (fibonaccichick666):

yupppers, just make sure you do your substitutions right

OpenStudy (fibonaccichick666):

hmm, not easily

OpenStudy (fibonaccichick666):

\[u=Cos^2(12x) ~~~~~~~~~~~~~~~~~~dv=Cos(12x)\] \[du=24cos12xsin12xdx~~~~~~~~~~~~v=1/12sin12x\] It be very messy and You'd end up with almost the exact same thing

OpenStudy (fibonaccichick666):

as that simple sub.

OpenStudy (anonymous):

Oh okay thank you very much again.

OpenStudy (fibonaccichick666):

np have a good one

OpenStudy (anonymous):

u 2 :)

OpenStudy (anonymous):

\[\int\limits \cos ^3(12x) dx=\frac{ 1 }{ 4 } \int\limits \left( \cos \left( 36x \right)+3\cos \left( 12x \right) \right)dx\] \[=\frac{ 1 }{ 4 }\left[ \frac{ \sin 36x }{36 }+3\frac{ \sin 12x }{12 } \right]+c\]

OpenStudy (fibonaccichick666):

Assuming the identity is right, that would work too^